Kevin ran 4 miles more than Steve ran. The sum of their distance is 26 miles.
How far did Steve run? The domain of the solution is {0, 4, 22, 11}.

Answer by nomad_55(6)

x=kevin
y=steve
x+y=26

x=4+y

(4+y)+y=26
-4 -4

y+y=22

2y=22
/2 /2

y=11

steve ran 11 miles.